### v2

# EdgeProfiler: A Fast Profiling Framework for Lightweight LLMs on Edge Using Analytical Model

Abstract
This paper introduces EdgeProfiler, a fast profiling framework designed for evaluating lightweight Large Language Models (LLMs) on edge systems. While LLMs offer remarkable capabilities in natural language understanding and generation, their high computational, memory, and power requirements often confine them to cloud environments. EdgeProfiler addresses these challenges by providing a systematic methodology for assessing LLM performance in resource-constrained edge settings. The framework profiles compact LLMs, including TinyLLaMA, Gemma3.1B, Llama3.2-1B, and DeepSeek-r1-1.5B, using aggressive quantization techniques and strict memory constraints. Analytical modeling is used to estimate latency, FLOPs, and energy consumption. The profiling reveals that 4-bit quantization reduces model memory usage by approximately 60–70%, while maintaining accuracy within 2–5% of full-precision baselines. Inference speeds are observed to improve by 2–3$\times$ compared to FP16 baselines across various edge devices. Power modeling estimates a 35–50% reduction in energy consumption for INT4 configurations, enabling practical deployment on hardware such as Raspberry Pi 4/5 and Jetson Orin Nano Super. Our findings emphasize the importance of efficient profiling tailored to lightweight LLMs in edge environments, balancing accuracy, energy efficiency, and computational feasibility.
Index Terms:
Large Language Models, LLM, TinyML, Edge Computing, Quantization, Low-Power Devices, Microcontrollers
I Introduction
Advances in Artificial Intelligence (AI) and deep learning have fueled remarkable progress in natural language processing (NLP), enabling applications such as summarization, text generation, question answering, and more. At the forefront of these developments are LLMs, which demonstrate an unprecedented ability to interpret and generate human-like text. LLMs have contributed to breakthroughs in mobile applications, healthcare, and situational analysis, among other domains. However, these models typically require substantial computational resources, memory, and power, which confine their deployment to cloud environments equipped with high-performance servers and GPUs.
Despite these challenges, there is growing interest in implementing LLMs on edge devices, including microcontrollers and low-power platforms. Edge deployment offers critical advantages in environments where cloud connectivity is unreliable or unavailable. For instance, LLMs can improve situational awareness during emergencies and disaster response, where cloud infrastructure may be compromised [1]. In these scenarios, devices running LLMs can provide real-time crisis management and communication support. Additionally, on-device processing enhances data security by minimizing exposure to vulnerabilities associated with transmitting data over the internet.
Reducing dependency on cloud connectivity is also a significant advantage. Moving LLM inference to the edge allows applications to function independently of network conditions, ensuring availability even in areas with limited or no internet access. For example, a local LLM could manage user notes or provide contextual assistance directly on a smartphone, regardless of network strength. Moreover, local inference significantly reduces latency, which is crucial for real-time applications. Unlike cloud-based models whose response times are affected by network stability and speed, on-device inference ensures prompt interaction, enhancing the user experience. Fig. 1 illustrates the process where LLM training is performed in a cloud GPU cluster within a data center environment, and pre-trained LLM models are then deployed to edge devices, which independently handle local inference computation.
However, deploying LLMs on edge devices remains challenging due to their complex architectures and high memory and power requirements [2]. Models such as GPT-3 and LLaMA, with billions of parameters, are not feasible for resource-constrained environments [3, 4]. This has led to an active area of research focused on quantization techniques, model pruning, and efficient inference strategies that adapt LLMs to the limitations of edge hardware [5]. These efforts aim to strike a balance between computational feasibility, energy efficiency, and model accuracy.
Figure 1: LLM inference deployment on diverse edge devices, highlighting the critical need for accurate profiling to ensure efficient computation, reduced latency, and optimized energy consumption in resource-constrained setups.
In this paper, we introduce EdgeProfiler, a fast profiling framework designed to systematically evaluate the performance of lightweight LLMs in resource-constrained environments. The findings indicate that aggressive quantization significantly reduces the memory and computational requirements of lightweight LLMs, enabling practical deployment on resource-constrained edge devices. The slight impact on model accuracy is outweighed by improvements in memory efficiency and faster inference speeds. Reduced energy consumption from low-bit precision enhances feasibility for power-limited environments, such as mobile devices and microcontrollers. Careful quantization strategies and efficient hardware utilization bridge the gap between LLM capabilities and the limitations of edge systems, paving the way for broader and more practical applications. The main contributions of this paper are as follows:
  • We introduce EdgeProfiler, a fast profiling framework for lightweight LLMs on edge devices.
  • Systematic evaluation of quantization and low-bit implementations to address memory and power constraints.
  • Analysis of strategies to enable LLM inference on microcontrollers and low-power edge platforms.
  • Comprehensive review of recent research and experimental results on deploying LLMs in hardware-constrained environments.
The structure of the paper is as follows: Section II presents a technical overview that highlights essential techniques applied to AI models for improving application-specific adaptations. Section III provides a detailed functional overview of the EdgeProfiler framework and the specific techniques employed in this study. Section IV discusses the evaluation results, while Section V offers a review of related work. Finally, Section VI summarizes the key insights and conclusions from the study.
II Background
Efficient deployment of LLMs on edge devices requires a deep understanding of techniques that balance computational cost, memory efficiency, and model accuracy. This section provides the foundational concepts and methodologies essential for such deployments. Key strategies include quantization approaches that reduce model size and computation while maintaining acceptable accuracy levels, and advanced techniques like Quantization-Aware Training (QAT) that ensure models remain robust under aggressive compression. By understanding the trade-offs between different quantization schemes and how they interact with model structure and hardware constraints, this background forms the basis for evaluating and optimizing LLM performance in resource-constrained environments.
II-A Quantization techniques for efficient LLM inference
Quantization refers to the process of reducing the numerical precision of model parameters and activations from high-precision FP16 / FP32 format to lower-precision representations of 8-bit integer or 4-bit integer[6].
Symmetric vs. Asymmetric. Quantization technique can be broadly classified into symmetric and asymmetric schemes, depending on how they map floating-point values to lower-precision integer representations.
II-A1 Symmetric Quantization
In symmetric quantization, zero in floating-point space is exactly representable by zero in integer space. The mapping function can be defined as:
$x_{\text{int}}=\text{round}\left(\frac{x_{\text{float}}}{s}\right)$ (1)
where $x_{\text{float}}$ is the original floating-point value, $x_{\text{int}}$ is the quantized integer value, and $s$ is a positive scaling factor (common across the tensor). The dequantization process simply scales back:
$x_{\text{float}}\approx s\times x_{\text{int}}$ (2)
II-A2 Asymmetric Quantization
In asymmetric quantization, a nonzero offset is introduced to better handle data distributions that are not centered at zero. The quantization function becomes:
$x_{\text{int}}=\text{round}\left(\frac{x_{\text{float}}-z}{s}\right)$ (3)
and the corresponding dequantization is:
$x_{\text{float}}\approx s\times x_{\text{int}}+z$ (4)
where $z$ is the zero-point, chosen such that zero in the integer domain corresponds to a nonzero floating-point value.
Asymmetric quantization is generally used for activations, where the dynamic range shifts during inference. Symmetric quantization is highly hardware-efficient because it eliminates the need for offset additions during inference. It is particularly well-suited for weight matrices, where the value distributions are typically centered around zero. An advantage of symmetric quantization is that it requires less memory and simpler hardware logic[7, 8]. However, a key disadvantage is that it can result in a higher mean squared error (MSE) compared to asymmetric quantization, especially when the data distribution is symmetric. Asymmetric quantization provides better representational fidelity for non-centered data but introduces extra computational overhead due to the zero-point handling.
Per-Tensor vs. Per-Channel Quantization. Quantization can be applied either globally across an entire tensor or separately across individual channels. These two strategies, per-tensor quantization and per-channel quantization, trade off between simplicity and representational accuracy[8].
II-A3 Per-Tensor Quantization
In per-tensor quantization, a single scale factor and, optionally, a single zero-point are used to quantize the entire tensor. The same quantization parameters are applied uniformly across all elements. similar to the equation 3. Per-tensor quantization offers the advantages of simple implementation and high efficiency, particularly for hardware accelerators that prefer uniform scaling across all elements. However, it also has notable disadvantages: it fails to capture the variance across different channels or feature maps, and it can lead to larger quantization errors if some channels have significantly wider or narrower value ranges compared to others[9].
II-A4 Per-Channel Quantization
In per-channel quantization, each channel of each output feature in a convolutional layer or each row in a matrix has its own scale and zero-point. This allows finer control over the quantization process.
$x_{\text{int},c}=\text{round}\left(\frac{x_{\text{float},c}-z_{c}}{s_{c}}\right)$ (5)
where $s_{c}$ and $z_{c}$ are the scale and zero-point specific to channel $c$. Per-tensor quantization is preferable when memory bandwidth and hardware simplicity are prioritized. In contrast, per-channel quantization provides better model accuracy, particularly for deep neural networks where channel-wise variation is significant. Modern LLM quantization schemes often apply per-channel quantization to weights and per-tensor quantization to activations, balancing accuracy and inference efficiency[10].
Figure 2: Quantization process from FP16 to INT8, demonstrating how a weight like 93.47 is scaled and rounded to 120 in INT8. This reduces memory size and computation while introducing minimal error, as the scaling factor preserves relative weight magnitudes.
Quantization-Aware Training (QAT). Quantization-Aware Training is a technique where quantization effects are simulated during the training phase itself, allowing the model to adapt to low-precision representations. During QAT, both the forward and backward passes emulate quantized operations, typically using fake quantization functions that maintain gradient flow[7]. This enables the model to learn parameters that are more robust to quantization noise, resulting in significantly better accuracy compared to naive post-training quantization[11]. Formally, during QAT, the model optimizes the following objective:
$\min_{\theta}\;\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\mathcal{L}\left(Q(f(x;\theta)),y\right)\right]$ (6)
where $f(x;\theta)$ denotes the model with parameters $\theta$, $Q(\cdot)$ denotes the quantization function, $\mathcal{L}(\cdot,\cdot)$ is the loss function, and $\mathcal{D}$ is the training data distribution.
QAT often uses lower bit-widths 8-bit or 4-bit during simulation while keeping high-precision master weights for gradient updates. By introducing quantization noise during training, QAT yields models that maintain higher accuracy after deployment, especially important for aggressive compression regimes in edge and embedded systems[3]. Despite promising benefits, LLM quantization presents unique challenges, such as attention layers and softmax output are particularly sensitive to quantization noise. Even if inputs are quantized to 4-bit/8-bit, intermediate results often require higher precision to avoid numerical instability. Some input sequences containing rare or domain-specific tokens can significantly stress low-precision models[10, 12]. Hardware supports efficient low-bit general matrix-matrix multiplication (GEMM) for arbitrary bitwidths, necessitating custom CUDA kernels. Fig. 2 shows the Symmetric quantization technique.
TABLE I: Comparison of specifications for Raspberry Pi 4, Raspberry Pi 5, and Jetson Orin Nano Super devices.
Device CPU RAM Storage OS
Raspberry Pi 4 Quad-core Cortex-A72 @1.5 GHz 2 GB/4 GB/8 GB LPDDR4 microSD card Linux
Raspberry Pi 5 Quad-core Cortex-A76 @2.4 GHz 4 GB/8 GB/16 GB LPDDR4X microSD card, PCIe storage Linux
Jetson Orin Nano Super 6-core Cortex-A78AE @1.7 GHz 8 GB 128-bit LPDDR5 microSD card, PCIe storage Linux
The paper [13] provides more insight into the numerous uses of LLM implementation in edge devices, in addition to quantization and compression strategies that makes this integration possible. These applications all emphasize the importance of utilizing LLMs locally for resource-constrained devices.
Figure 3: High-level overview of the EdgeProfiler. The framework integrates model, hardware, and precision parameters to estimate latency, memory footprint, and energy consumption, enabling exploration of performance trade-offs on edge platforms.
III Methodology
Fig. 3 presents a high-level overview of our lightweight profiling framework, which evaluates the end-to-end performance and energy cost of serving transformer-based language models on a range of edge platforms. The framework operates through a modular workflow: it takes as input the model configuration, hardware specifications, and numeric precision settings; internally, it performs analytical modeling of key performance metrics such as parameter count, FLOPs, memory footprint, latency, and energy consumption; and finally, it outputs detailed performance estimates, enabling trade-off exploration across different device-model-precision combinations. This setup provides a scalable method for understanding and comparing LLM behavior on resource-constrained systems without requiring full hardware deployment.
III-A Performance Modeling Framework
At the core of our framework is the EdgeProfile, which takes three inputs: model configuration, hardware configuration, and precision configuration.
Model configuration. Specifies transformer architecture parameters: number of layers ($L$), hidden dimension ($H$), intermediate dimension ($I$), attention heads, vocabulary size ($V$), and sequence length ($S$).
Hardware configuration. Encapsulates peak computational throughput (FLOPs/sec), DRAM bandwidth, storage throughput, PCIe host‐to‐device bandwidth, interconnect/network bandwidth, utilization factors ($U_{\text{compute}}$, $U_{\text{memory}}$, $U_{\text{storage}}$, $U_{\text{H2D}}$, $U_{\text{net}}$), and energy cost per flop or per byte accessed.
Precision configuration. Defines the data‐type size $B$ in bytes (4B for FP32, 2B for FP16, 1B for INT8).
EdgeProfiler computes:
Parameter count
$P=L\,(4H^{2})+L\,(2HI)+2VH$ (7)
Equation 7 captures all of the model’s weight parameters. The four projection matrices per layer plus the two feed-forward matrices, and the input/output embeddings.
FLOPs per token
$\mathrm{FLOPs/token}$ $=L\bigl{(}6H^{2}+4HS+4HI+4IH+9H\bigr{)}$ (8)
Equation 8 accounts for both attention projections query, key, value, and output projection(Q,K,V,O) and the two matmuls in each feed‐forward block, as well as LayerNorm and Softmax overhead.
Memory footprint
$M=P\cdot B+S\cdot H\cdot B+2L\cdot S\cdot H\cdot B$ (9)
Equation 9 includes storing all weights ($P\cdot B$), the hidden‐state activations for the full sequence ($S\!H\!B$), and the cached key/value pairs across layers ($2L\!S\!H\!B$).
III-B Latency Breakdown
Using the counts above, we estimate each component of on-chip inference time.
Compute latency
$T_{\mathrm{comp}}=\frac{\mathrm{FLOPs/token}}{\mathrm{peak}_{\mathrm{flops}}\times U_{\mathrm{compute}}}$ (10)
Equation 10 performs all the arithmetic at the device’s effective compute throughput.
Memory‐bound latency
$T_{\mathrm{mem}}=\frac{M}{\mathrm{memory}_{\mathrm{BW}}\times U_{\mathrm{memory}}}$ (11)
Equation 11 captures the delay reading/writing weights, activations, and cache from on-chip memory.
I/O, Host‐to‐Device, and Network Overheads
Each is computed by dividing the total weight bytes ($P\cdot B$) or shard bytes by the respective bandwidth and utilization factor:
$T_{\mathrm{I/O}}=\frac{P\,B}{\mathrm{STORAGE}_{\mathrm{BW}}\times U_{\mathrm{storage}}}$ (12)
$T_{\mathrm{h2d}}=\frac{P\,B}{\mathrm{H2D}_{\mathrm{BW}}\times U_{\mathrm{H2D}}}$ (13)
$T_{\mathrm{net}}=\frac{S\,H\,B}{\mathrm{NET}_{\mathrm{BW}}\times U_{\mathrm{net}}}$ (14)
Equation 12 is the latency to load model weights from disk or flash storage. Equation 13 is the latency to copy weights over PCIe or NVLink into GPU memory. Equation 14 is the latency to exchange a shard of key/value cache across nodes. We also provide a fine‐grained breakdown of attention‐projection, key/value matmuls, MLP, layernorm, and softmax by dividing each operator’s raw FLOP count by effective compute throughput.
III-C Energy Modeling
Energy per token is estimated as the sum of compute and data‐movement costs:
$E=\mathrm{FLOPs/token}\,\times\,e_{\mathrm{flop}}\;+\;M\,\times\,e_{\mathrm{byte}},$ (15)
where $e_{\mathrm{flop}}$ and $e_{\mathrm{byte}}$ are measured in joules per FLOP and per byte, respectively.
This comprehensive modeling approach allows for systematic exploration of design trade-offs, helping identify bottlenecks and optimize configurations for efficient LLM deployment on edge systems.
Figure 4: Latency and energy profile using EdgeProfiler: (a) Memory-bound latency, (b) Storage I/O latency to load weight from disk, (c) Host-to-device transfer time, (d) Network-bound latency for a single KV-shard exchange across nodes, (e) End-to-end latency per token, (f) Estimated energy per token.
TABLE II: Comparison of model size, memory usage, inference speed, and accuracy loss across different quantization precisions.
Model Precision Model Size Memory at Runtime Inference Speed Accuracy Loss
FP16 2.2GB $\sim$3.13GB 1$\times$ baseline
TinyLlama INT8 1.2GB $\sim$2.25GB 1.86$\times$ minor
INT4 644MB $\sim$1.78GB 2.45$\times$ moderate
FP16 2.0GB $\sim$2.44GB 1$\times$ baseline
Gemma3-1B INT8 1.1GB $\sim$1.60GB 1.26$\times$ minor
INT4 815MB $\sim$1.35GB 1.52$\times$ moderate
FP16 2.5GB $\sim$3.58GB 1$\times$ baseline
Llama3.2-1B INT8 1.3GB $\sim$2.53GB 2.7$\times$ minor
INT4 776MB $\sim$2.01GB 3.33$\times$ moderate
FP16 3.6GB $\sim$3.91GB 1$\times$ baseline
DeepSeek-r1-1.5B INT8 1.9GB $\sim$2.55GB 2.19$\times$ minor
INT4 1.1GB $\sim$1.84GB 2.97$\times$ moderate
IV Evaluation
This section evaluates the performance of lightweight LLMs on representative edge devices. The profiling framework is used to analyze trade-offs between model size, precision levels, and hardware characteristics. Key metrics such as latency, memory footprint, arithmetic intensity, and energy consumption are examined, offering insights into the feasibility and efficiency of deploying LLMs on resource-constrained platforms. The framework was executed on a standard workstation equipped with a 10th Gen Intel(R) Core(TM) i7-10700F CPU, 32 GB DDR4 memory, and without using a dedicated GPU, running Ubuntu 22.04 LTS.EdgeProfiler is available on GitHub: EdgeProfiler
Experimental Setup. We instantiate the profiler on three edge platforms: Raspberry Pi 4, Raspberry Pi 5, and Jetson Nano Super, using published peak FLOPs and bandwidths with calibrated utilization factors. The analysis covers multiple numeric precisions (FP32, FP16, INT8) and representative 1–1.5B parameter LLMs. For each configuration, the profiler outputs key metrics, including parameter count, FLOPs per token, peak memory footprint, stage-wise latency (compute, memory, I/O, H2D, network), end-to-end latency, arithmetic intensity, and energy per token. This analytical approach enables rapid comparison of architecture and precision trade-offs without requiring full hardware deployment. Details of the hardware configurations are summarized in Table I.
Profiling Results Analysis. On all three devices, storage I/O accounts for the vast majority of end-to-end latency shown in Fig. 4(b). Even though compute (and memory) times are on the order of a few hundred milliseconds or less, I/O delays range from multiple seconds (Raspberry Pi 4/5) down to just under a second on the Jetson Nano Super. This indicates that, without specialized weight-loading optimizations, simply reducing arithmetic cost (e.g., via quantization) will yield diminishing returns once compute time becomes negligible relative to data-movement overhead.
Precision reduction from FP32 to FP16 halves each component’s latency, and INT8 cuts it roughly by four. On Raspberry Pi 4, end-to-end latency drops from $\sim$15.4 s (FP32) to $\sim$3.9s (INT8), driven almost entirely by shorter I/O and transfer times of smaller weight footprints. However, even at INT8, I/O remains the bottleneck (3.5s vs. compute  0.13s), underscoring that the network and storage subsystems must be improved in tandem with quantization to realize truly low-latency inference shown in all the Fig. 4.
The Jetson Nano Super’s higher storage bandwidth and PCIe host-to-device throughput, shown in Fig. 4(c), deliver a dramatic reduction in I/O cost: INT8 inference completes in $\sim$1.05s end-to-end, nearly four times faster than on the Raspberry Pi 5. Compute and memory latencies on the Jetson ($\sim$0.07s and $\sim$0.88s at FP32) are comparable to—though still lower than those on the Pis, but the real advantage comes from overlapping and hiding I/O behind faster transfers. This highlights that mid-range AI accelerators can shift the bottleneck away from storage if paired with efficient weight-delivery mechanisms.
Across all models and platforms, arithmetic intensity remains low (well under 1 FLOP/byte), confirming a data-movement–bound regime. As models grow, compute time will rise, but so too will weight volume, so similar bottlenecks are likely unless novel compression or on-device caching techniques reduce the need for repeated weight fetches. In practice, combining aggressive quantization with persistent in-RAM weight formats, alongside higher-bandwidth storage paths (e.g., NVMe, LPDDR5X), appears essential for sub-second LLM inference at the billion-parameter scale.
Ablation Studies on Quantization. Table II summarizes how reducing precision from FP16 to INT8 and INT4 impacts model size, peak memory usage, inference throughput, and observed accuracy degradation across four 1-1.5B (TinyLlama‐1B, Gemma3‐1B, Llama3.2‐1B, DeepSeek‐r1‐1.5B) and peak runtime memory (3.13GB to 2.25GB), while boosting inference speed by roughly 1.8×1.8$\times$1.8 × with only a minor accuracy loss. Further quantization to INT4 yields even more dramatic storage saving (down to 644MB for TinyLlama) and a 2.45×2.45$\times$2.45 × speedup, at thecost of a moderate drop in task performance,
Across all architectures, INT8 delivers near-2×2$\times$2 × speed gain and $\sim$50reduction in memory footprint with negligible impact on accuracy. The large models Llama3.2-1B and DeepSeek-r1-1.5B benefit even more from low-precision storage, up to $\times3$ speedup, incur more noticeable accuracy degradation, suggesting their suitability only when memory and latency constraints are extreme and a moderate loss in quality is acceptable.
V Related Works
Recent research has focused on adapting LLMs for edge devices, addressing challenges like limited memory, compute capacity, and energy constraints. Work [14] empirically studies how resource-constrained computing environments affect the design choices for personalized LLMs, uncovering key trade-offs and offering practical guidelines for customization and deployment on edge devices. However, the empirical study might be limited to specific models and hardware, affecting generalizability. Deeploy [15] introduces a compiler framework that translates pre-trained Foundation Models into compact Small Language Models (SLMs) for microcontrollers, showcasing a practical approach for edge NLP tasks. Post-training quantization techniques, such as GPTQ [16], compress model weights to as low as 3–4 bits, balancing memory efficiency and accuracy. Complementary methods like AWQ [17] and SmoothQuant [18] further enhance memory and performance trade-offs.
MobileLLM [19] demonstrates that sub-billion parameter models can provide effective NLP functionality on mobile and edge platforms, achieving a balance between resource constraints and utility. BioMistral-7B [20] applies quantization and model merging to tailor LLMs for biomedical tasks, combining compression with domain-specific optimization for constrained environments. Memory innovations, exemplified by “LLM in a Flash” [21], reduce DRAM usage by dynamically loading only essential model weights from flash storage during inference. This approach addresses limitations of static weight loading, improving LLM feasibility for edge deployment. Another recent advancement in efficient LLM deployment is the work on automatic INT4 weight-only quantization and optimized CPU inference, as described by the authors in [22]. This approach introduces an automatic quantization flow and a specialized LLM runtime with highly optimized kernels, enabling extreme inference efficiency on CPUs. Their method is demonstrated on popular models including LLaMA2, LLaMA, and GPT-NeoX, highlighting the potential for practical, low-precision LLM deployment on standard hardware. This complements efforts like GPTQ and “LLM in a Flash” by extending efficient deployment strategies to CPUs, further broadening the applicability of LLMs in resource-constrained environments.
The works discussed collectively highlight various strategies such as compact model generation, quantization, and innovative memory management are necessary to deploy LLMs on resource-constrained edge devices. This underscores the pressing need for a comprehensive profiling framework specifically designed for edge environments. To the best of our knowledge, no existing approach offers an analytical modeling solution for this challenge. EdgeProfiler addresses this gap by taking both the model and hardware configuration as inputs, while targeting a desired level of precision in the performance evaluation. It does not involve architecture level simulators such as gem5 [23] rather a set of analytical models which makes it faster profiling and is easier to implement.
VI Conclusion
In this paper we have propose EdgeProfiler, a fast profiling framework for lightweight LLMs on edge devices using an analytical model. EdgeProfiler framework reveals how hardware characteristics, model size, and numerical precision jointly shape inference performance and efficiency. We present our findings from three widely used edge devices. Across all platforms, reducing precision from FP32, FP16, and INT8 yields substantial end-to-end latency and energy savings. INT8 cuts the latency by $\sim$75and energy by $\sim$75relative to FP32. On low-power devices, Raspberry Pi, memory, and I/O bound stage often dominate, whereas on more capable hardware like Jetson Nano, compute latency becomes negligible, and I/O or network transfer surface as the new bottleneck. Building upon this foundation, future work will focus on refining profiling granularity, integrating real-world deployment feedback, and expanding support for emerging hardware platforms to accelerate practical, scalable, and energy-efficient LLM applications on the edge.